OpenStudy (anonymous):

If cos θ = 8/17 and 3π/2 < θ < 2π, compute sin θ and cot θ. 1) sin θ= 2) cot θ=

OpenStudy (jhannybean):



OpenStudy (jhannybean):

\[\sf \text{SOHCAHTOA} \implies \cos(\theta) =\frac{adj}{hyp}\]

OpenStudy (jhannybean):

therefore \(\sf \sin(\theta) = \dfrac{opp}{hyp}\)

OpenStudy (jhannybean):

We can figure out the opposite side by using pythagorean theorem: \(y^2 = 17^2-8^2\)

OpenStudy (jhannybean):

\[y= \sqrt{17^2-8^2}\]

OpenStudy (jhannybean):

So therefore, \(\sin(\theta) = \frac{y}{17}\)

OpenStudy (anonymous):

It has to be an exact answer, it won't accept square roots.

OpenStudy (jhannybean):

You didn't even solve for it.

OpenStudy (anonymous):

the square root IS the exact answer

OpenStudy (jhannybean):

gotta go.

OpenStudy (anonymous):

Oh wait, I got it, thank you.